how do u solve this problem and what would be the point i put on the graph... the questions goes.
Samantha earns $7 per month more than her sister. write an equation to find the amount of money, y, samantha earns, when her sister ears x dollars. graph the equation in the space at the right... the it gives me some space to draw the graph i just need to know how to solve the problem and what the points (x,y) are gonna be....



then it tells me to use the graph that im going to supposebly draw to find how much samantha earns when her sisters earns $2.50. How can y ou check your answer? (i think that doesn't make sence and thats why i need help on learning how to solve this.........

Answer by stanbon(48551)

Let amount her sister earns be "x"
The samantha earns "x+7" dollars.
EQUATION:
y=x+7
Plot two points:
If x=0 y=0+7=7 gives the point (0,7)
If x=5 y=5+7=12 gives the point (5,12)
Plot those and draw a line through them carefully.
Now see what the "y" value would be if x were 2.5.
You can check your estimated answer by putting 2.5
into the equation in place of x:
y=2.5+7=9.5
Cheers,
Stan H.